# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

### v1

@@ 3 Hanabi
The Hanabi problem along with how human conventions are used to solve the problem effectively, must first be discussed in order to give context for our novel approach. Hanabi is a 2–5 player card game, best described as a cooperative solitaire [36]. Players have a hand of five cards for two and three player, and four cards for four and five player. Each card has a suit/colour (red, yellow, green, white, or blue) and a rank (1 to 5). The deck comprises 50 cards total, with 10 cards for each suit with a distribution of three 1’s, two 2’s, 3’s, and 4’s, and finally only one 5. The aim of the game is to stack these suits in numerical order from 1 to 5, however players cannot see their own cards, only the cards of their fellow players.
@@
Figure 1: An example game of Hanabi as seen from the perspective of player 1. It is now player 1’s turn, and they can take the hint, play or discard action.
At the start of the game the players have a shared total of three life tokens and eight hint tokens. If all the hint tokens are depleted, a player cannot take the hint action and must either play or discard a card from their hand. The players are awarded a shared score depending on the number of cards successfully placed on each stack. If the players manage to build the stacks to the maximum of 5 each, the game ends and the players receive a perfect score of 25/25. Alternatively, if the players lose all of their life tokens (called “bombing out”), or if the deck has been depleted, the game also ends. Note, bombing out results in a score of 0/25 independent of how many cards were played successfully. A discard action, or a misplay, will result in the card being removed from the game, and that player must draw a new card from the deck to replenish the missing card (which is also hidden from current the player).
@@ 3.1 Human Conventions in Hanabi
Due to the nature of Hanabi, players have too few hint tokens in order to effectively convey all the needed information (colour and rank) of the 25 playable cards [36]. This becomes even more challenging as the number of players increase and hint tokens become more valuable, as well as the number of playable turns decreasing [^5] As the number of players increase, the cards remaining in the deck decrease, and a single player has less playable turns.. Hints must therefore be used wisely, and must convey information about more than just one card. This is possible since more than one card with a similar rank or suit can be touched (or revealed) by a single hint, along with negative information [^6] Negative information refers to implied information about non-focus cards, e.g. if a player hints that two cards are green, the other cards must therefore be non-green. being given about all the other cards in a hand. Unfortunately, this is not enough, and players who don’t use additional strategies will struggle to reliably achieve a score above 15/25.
@@
Players, therefore, require a way to convey additional information through implicit communication, while still remaining within the rules of the game. As stated by Bard et al. [36]: “This implicit information is not conveyed through the impact that an action has on the environment (i.e., what happens) but through the very fact that another player decided to take this action (i.e., why it happened)”. Players can then reason over the actions of others, and implicitly communicate through them. This is done through the use of conventions, or mutually agreed upon “rules”, that players develop outside the game. For example, in Fig. 1 the players have a convention where an ambiguous hint has the focus placed on their left-most card, therefore player 1 can hint to player 2 that they have two red cards in position one and three respectively, and player 2 will know (as a result of their convention) that the left-most card must be the playable red 1. Since the game’s release, players have developed their own intricate and extensive conventions, often specific to their group of friends or cooperators, examples include the Board Game Arena [^7] https://forum.boardgamearena.com/viewtopic.php?t=5252 and the H-Group [^8] https://hanabi.github.io/, with the H-Group being the most widely used and adapted.
4 Augmenting the Action Space with Conventions
Existing MARL approaches specific to Hanabi naturally develop conventions, but these conventions are nonsensical and vary greatly from run to run [27, 47, 48]. We propose incorporating existing human conventions into MARL algorithms using artificial conventions, which act as advanced cooperative actions that span over multiple time steps and agents. Conventions require that all the agents involved participate or “subscribe” in order for it to reach fruition, however the agents cannot communicate directly which convention is being started or followed.
Conventions can be divided into two categories, namely available conventions and active conventions. We define a convention $C$ as
$C=\langle I,\pi_{I},M,\pi_{M},F,\pi_{F}\rangle,$ (5)
where the initial condition $I$ must be met for a convention to become available ($O_{t}\in I$), and if an agent chooses to initiate $C$, the policy $\pi_{I}$ determines the corresponding environment action to start the convention. The convention will then become active, and another agent can “subscribe” to $C$ if they meet the condition $M$, with $\pi_{M}$ determining their environment action to continue $C$. Eventually an agent will have the opportunity to perform a unique continuation action to complete $C$, but only if they meet the final condition $F$, with $\pi_{F}$ determining their environment action to terminate $C$ and allow it to reach completion. For the sake of simplifying this discussion, and to be consistent with our implemented Hanabi conventions in later sections, we will focus on two-step conventions, where one agent initiates a convention $C$ and another agent must opt in to complete the convention, i.e. $C=\langle I,\pi_{I},F,\pi_{F}\rangle$. The concept behind conventions is best described at the hand of an example:
@@ 4 Augmenting the Action Space with Conventions
Conventions Example. In Fig. 1 each player is controlled by an agent, with the current player (referred to as the active player) being agent 1. The convention at hand, which we will refer to as $c_{1}$, states that if an ambiguous hint is given, the focus of that hint is the left-most (or newest) card [^9] This is a common human convention and is based on the ”Single card focus” principle as discussed by the H-Group.. Since the condition $I_{c1}$ is met, i.e. player 2 has two cards in their hand that share the same colour or rank and the left-most one is playable, the action {start $c_{1}$} is available for agent 1. Agent 1 chooses to initiate $c_{1}$ which in turn gets translated by the policy $\pi_{Ic1}$ to an environment action of {hint red to player 2}, ending player 1’s turn. It is now player 2’s turn, and they have just received an ambiguous colour hint from player 1, and given their current observation they can derive that $c_{1}$ is currently active. This triggers the subscribing (and also completing) condition $F_{c1}$, where agent 2 can now opt in to continue (and complete) $c_{1}$. When agent 2 chooses to subscribe to $c_{1}$, their action is translated by the policy $\pi_{Fc1}$ to an environment action of {play card in position 1}, subsequently ending player 2’s turn and completing $c_{1}$.
It is important to note that conventions do not have to be sequential, e.g. in the case of the example above, if agent 3 had their turn in between agent 1 and 2, agent 3 would be able to initiate, or subscribe to, a different convention (since they do not meet any conditions of $c_{1}$) and $c_{1}$ will remain active. When more conventions are introduced, the learning problem becomes apparent and crucial, since agents must learn which convention is applicable in certain scenarios where more than one convention is available. Furthermore, since agents have the ability to opt in to a convention, they can also choose not to, which could result in an active convention termination without a completing action. This allows agents to halt certain conventions based on their unique observations alluding to it being non-optimal. Thus, the agents must learn when to prioritise superior conventions over existing or active inferior ones.
4.1 Architecture Design and Implementation
Algorithm 1 can be adapted to include conventions, and produce Algorithm 2, which serves as the pseudocode for conventions applied to independent Q-learning in a turn-based scenario. The current observation, which also contains the environment actions of the last round, determines the available and/or active conventions, shown in lines 9-15. Once all the conventions are determined, agent $i$ can choose to start or continue a convention based on the policy derived from $Q^{i}$ and their observation $O^{i}_{t}$, shown in line 16. Similar to human-play, the agents aren’t allowed to communicate which convention is currently being started or followed, but rather only observes the environment action determined by the appropriate convention policy, shown in line 22.
Algorithm 2 Independent Q-learning with conventions in a turn-based environment
1:Initialise hyperparameters similar to Algorithm 1
2:Define the list of conventions $\boldsymbol{C}$ with $C=\langle I,\pi_{I},F,\pi_{F}\rangle$
3:Initialise all $Q(O,C)\leftarrow 0$
4:for each episode do
5:     Reset environment and set total time $t\leftarrow 0$
6:     while $S_{t+1}$ is not terminal do
7:         for all players $i$ from 00 to $P-1$ do
8:              Obtain $O^{i}_{t}$ from environment
9:              for each convention $c$ in $\boldsymbol{C}$ do
10:                  if $O^{i}_{t}\in I_{c}$ then
11:                       Convention $c$ is available
12:                  else if $O^{i}_{t}\in F_{c}$ then
13:                       Convention $c$ is active
14:                  end if
15:              end for
16:              $C^{i}_{t}=\text{argmax}_{c}Q^{i}(O^{i}_{t},c)$triangleright Choose $C^{i}_{t}$ from the available and/or active conventions $\boldsymbol{c}$ based on $O^{i}_{t}$ and the policy derived from $Q^{i}$
17:              if $C^{i}_{t}$ was available then
18:                  $A^{i}_{t}=\pi_{Ic}(O^{i}_{t})$triangleright $\pi_{Ic}$ determines the environment action $A^{i}_{t}$
19:              else if $C^{i}_{t}$ was active then
20:                  $A^{i}_{t}=\pi_{Fc}(O^{i}_{t})$triangleright $\pi_{Fc}$ determines the environment action $A^{i}_{t}$
21:              end if
22:              Take action $A^{i}_{t}$ in the environment
23:              Receive and store $R_{t+1}$ and $O^{i}_{t+1}$, with $O^{i}_{t+1}\in S_{t+1}$, from environment
24:              $\sum_{k=0}^{P-1}R_{t+1+k}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
25:              $Q^{i}(O^{i}_{t},C^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},C^{i}_{t})+\alpha\bigr{[}\sum_{k=0}^{P-1}R_{t+1+k}+\gamma\max_{c}Q^{i}(O^{i}_{t+P},c)-Q^{i}(O^{i}_{t},C^{i}_{t})\bigr{]}$
26:              $t\leftarrow t+1$
27:         end for
28:     end while
29:end for
From the environment’s perspective agents are merely choosing actions based on the observations the environment provided, but in reality the agents are choosing these actions based on predefined convention policies $\pi_{I}$ or $\pi_{F}$. In the case of deep RL, lines 9–14 and 17–21 indicate that our approach requires two new layers to be added to the agent’s architecture, namely one before the input layer to determine the available and/or active conventions and one at the output to translate the chosen convention into an environment action. This is illustrated in Fig. 2, and in the case of deep Q-learning, the action selection step is performed using $\text{argmax}_{c}Q(O_{t},c)$, with the available and/or active conventions acting as an action mask.
Figure 2: Architecture design for a feed forward neural network (NN) with conventions applied. The environment provides the observation $O_{t}$ and receives the action $A_{t}$. The size of the input layer is equal to the size of the observation tuple ($x$), and each hidden layer has a size equal to the number of atoms for that layer. The size of the output layer is equal to twice the number of conventions ($2y$), since we focus on two-step conventions, and the convention selection is determined by the MARL algorithm.
4.2 Action Space Augmentation
However, there can exist scenarios where no conventions are available, and the agents will not be able to take any actions. To solve this problem, we propose augmenting the action space with primitive actions and conventions simultaneously, similar to the discussion by Sutton et al. [19] on combining primitive actions and options. This requires that the conventions be appended to the existing primitive-action space (with size $z$), and results in an action-convention space consisting of
$C^{\prime}=\{a_{0},a_{1},\dots,a_{z},c_{1},c_{2},\dots,c_{y}\}.$ (6)
We can adapt Algorithm 2 to include an augmented action-convention space, and produce Algorithm 3.
1:Initialise hyperparameters similar to Algorithm 1
2:Define the list of convention $\boldsymbol{C}$ with $C=\langle I,\pi_{I},F,\pi_{F}\rangle$
3:Combine primitive actions and conventions to produce the augmented action-convention space $C^{\prime}$
4:Initialise all $Q(O,C^{\prime})\leftarrow 0$
5:for each episode do
6:     Reset environment and set total time $t\leftarrow 0$
7:     while $S_{t+1}$ is not terminal do
8:         for all players $i$ from 00 to $P-1$ do
9:              Obtain $O^{i}_{t}$ from environment
10:              for each convention $c$ in $\boldsymbol{C}$ do
11:                  if $O^{i}_{t}\in I_{c}$ then
12:                       Convention $c$ is available
13:                  else if $O^{i}_{t}\in F_{c}$ then
14:                       Convention $c$ is active
15:                  end if
16:              end for
17:              ${C^{i}_{t}}^{\prime}=\text{argmax}_{ac}Q^{i}(O^{i}_{t},ac)$triangleright Choose ${C^{i}_{t}}^{\prime}$ from the available and/or active conventions $\boldsymbol{c}$ or primitive actions $A$ based on $O^{i}_{t}$ and the policy derived from $Q^{i}$
18:              if ${C^{i}_{t}}^{\prime}\in\boldsymbol{C}$ then $\triangleright$ If the chosen action is a convention
19:                  if $C^{i}_{t}$ was available then
20:                       $A^{i}_{t}=\pi_{Ic}(O^{i}_{t})$triangleright $\pi_{Ic}$ determines the environment action $A^{i}_{t}$
21:                  else if $C^{i}_{t}$ was active then
22:                       $A^{i}_{t}=\pi_{Fc}(O^{i}_{t})$triangleright $\pi_{Fc}$ determines the environment action $A^{i}_{t}$
23:                  end if
24:              else if ${C^{i}_{t}}^{\prime}\in A$ then $\triangleright$ If the chosen action is a primitive action
25:                  $A^{i}_{t}={C^{i}_{t}}^{\prime}$
26:              end if
27:              Take action $A^{i}_{t}$ in the environment
28:              Receive and store $R_{t+1}$ and $O^{i}_{t+1}$, with $O^{i}_{t+1}\in S_{t+1}$, from environment
29:              $\sum_{k=0}^{P-1}R_{t+1+k}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
30:              $Q^{i}(O^{i}_{t},C^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},C^{i}_{t})+\alpha\bigr{[}\sum_{k=0}^{P-1}R_{t+1+k}+\gamma\max_{ac}Q^{i}(O^{i}_{t+P},ac)-Q^{i}(O^{i}_{t},C^{i}_{t})\bigr{]}$
31:              $t\leftarrow t+1$
32:         end for
33:     end while
34:end for
We can also adapt Fig. 2 to account for an augmented action-convention space by simply extending the output layer to include the primitive actions along with the conventions, and account for the difference between primitive action selection over convention selection in the convention-action translation. To explore and validate the capabilities of conventions and action space augmentation, we conduct preliminary tests using a simplified version of Hanabi, called Small Hanabi.

### v2

@@ 2.3 Options
3 Augmenting the Action Space with Conventions
Similar to how options enabled the incorporation of existing domain knowledge into RL algorithms by modifying an agent’s action space, we propose a MARL solution strategy to achieve a similar goal. Options required the defining of semi-MDPs, where actions can have varying lengths, and in a MARL scenario will theoretically require decentralised partially observable semi-MDPs (encapsulated by Amato et al.’s MacDec-POMDP [46]). However, our solution strategy does not directly change the length of an action by having an external policy take over, rather it facilitates implicit communication through a specific behaviour as a result of a sequence of actions, based on predefined principles, called a convention. Each convention contains convention-steps that span over and include multiple time steps and multiple agents, requiring them to actively opt in for the convention to reach fruition. Therefore, our augmentation of the action space does not directly change the Dec-POMDP framework, and merely acts as an extension of the existing action space to allow for more advanced behaviours.
Ideally an agent’s action space can be comprised of a pure-convention space, i.e. containing only convention-steps and no primitive actions. However, there can exist scenarios where no convention-steps are available, and the agents will not be able to take any actions. To solve this problem, we propose augmenting the action space with primitive actions and conventions simultaneously, similar to the discussion by Sutton et al. [19] on combining primitive actions and options at the slight cost of performance. We will prove in Section 4.3 at the hands of the Small Hanabi problem, that this solution strategy only slightly impacts the overall performance when comparing agents with a pure-convention space and an augmented action-convention space.
3.1 Conventions
Existing MARL approaches in partially observable environments, such as Hanabi, naturally develop conventions, but these conventions are nonsensical and vary greatly from run to run [27, 48, 49]. We propose incorporating existing human conventions into MARL algorithms using artificial conventions, which act as advanced cooperative actions that span over and include multiple time steps and agents. Conventions require that all the agents involved participate or “subscribe” in order for it to reach fruition, however the agents cannot communicate directly which convention is being started or followed.
Conventions can be divided into two categories, namely available conventions and active conventions, based on if an agent can initiate the convention or continue the convention. Each convention contains a number of steps for that convention (referred to as convention-steps) which corresponds to the number of actions in that convention, allowing it to reach fruition and result in a desired behaviour. We define a convention $c_{k}$, containing a number of convention-steps $m_{k}$, as
$c_{k}=\langle m_{k},\lambda_{k}^{1},\pi_{k}^{1},\lambda_{k}^{2},\pi_{k}^{2},\dots,\lambda_{k}^{m_{k}},\pi_{k}^{m_{k}}\rangle.$ (5)
The set of conditions $\lambda_{k}$ determine which convention-steps are currently available to an agent based on their current observation, i.e. where $O_{t}\in\lambda_{k}$ determines the available convention-steps $1:m_{k}$. Subsequently, the set of policies $\pi_{k}$ determine the corresponding environment action $A_{t}$ for each convention-step with $A_{t}=\pi_{k}(O_{t})$.
Thus, the initial condition $\lambda_{k}^{1}$ must be met for a convention to become available, and if an agent chooses to initiate $c_{k}$, the policy $\pi_{k}^{1}$ determines the corresponding environment action to start the convention. The convention will then become active, and agents can “subscribe” to $c_{k}$ if they meet any of the conditions $\lambda_{k}^{2:m_{k}}$, with $\pi_{k}^{2:m_{k}}$ determining their environment action to continue $c_{k}$. Eventually an agent will have the opportunity to perform a unique continuation action to complete $c_{k}$, but only if they meet the final condition $\lambda_{k}^{m_{k}}$, with $\pi_{k}^{m_{k}}$ determining their environment action to terminate $c_{k}$ and allow it to reach fruition. When implementing $K$ conventions in an agent’s action space, the convention-steps of the different conventions are combined to form the convention-step space [^6] Note that the number of implemented conventions $K$ is not equal to the size of the convention-step space, since conventions are variable length vectors based on the number of steps $m_{k}$ contained within each convention $c_{k}$. defined as
$C=\langle\pi_{0}^{1},\pi_{0}^{2},\dots,\pi_{0}^{m_{1}},\pi_{1}^{0},\pi_{1}^{1},\dots,\pi_{K}^{m_{K}}\rangle,$ (6)
with
$|C|=\sum_{j=0}^{K}m_{j}.$ (7)
Conventions are further distinguished based on their step-size $m_{k}$, with single-step conventions having a step-size of $m_{k}=1$, two-step conventions having $m_{k}=2$, and multistep conventions having $m_{k}>2$. Single-step conventions, i.e. $c_{k}=\langle 1,\lambda_{k}^{1},\pi_{k}^{1}\rangle$, contain only one condition $\lambda_{k}^{1}$, as well as one policy $\pi_{k}^{1}$, which simultaneously acts as the initial-and final conditions of the convention, allowing an agent to initiate and terminate the convention on a single time step. Although this might seem similar to a primitive action, single-step conventions still allow for additional communication through the observed behaviour resulting from the policy $\pi_{k}^{1}$, since the policy will result in a different environment action based on the current observation, as opposed to primitive actions which remain constant independent of the current observation. Two-step conventions only contain two conditions and two policies, which act as the initial-and terminating steps of the convention, i.e. $c_{k}=\langle 2,\lambda_{k}^{1},\pi_{k}^{1},\lambda_{k}^{2},\pi_{k}^{2}\rangle$.
It is important to note that conventions do not have to be sequential, e.g. if agent 3 had their turn in between agent 1 and 2, and convention $c_{1}$ is active and only relevant to player 1 and 2, agent 3 would be able to initiate or subscribe to a different convention and $c_{1}$ will remain active. When more conventions are introduced, the learning problem becomes apparent and crucial, since agents must learn which convention is optimal in certain scenarios where more than one convention is available. Furthermore, since agents have the ability to opt in to a convention, they can also choose not to, which will result in an active convention terminating without a completing action. This allows agents to halt certain conventions based on their unique observations alluding to it being non-optimal. Thus, the agents must learn when to prioritise superior conventions over existing or active inferior ones. We note that the majority of our implemented conventions, presented and discussed in Appendix A, are single-and two-step conventions, with the only multistep conventions being the Prompt and the Finesse (each having an $m_{k}=3$).
3.2 Action Space Augmentation
Conventions can either act as the only actions available to an agent, or can be incorporated into the existing action space using action space augmentation. This requires that the convention-step space be appended to the existing primitive-action space, and produce the augmented action-convention space. However, before this can be achieved, the primitive actions must be translated into unique conventions to ensure effective cohesion. Ultimately, a primitive action is a special type of single-step convention with no initial condition and a fixed deterministic policy mapping to a certain environment action independent of the current observation, i.e.
$c_{k}^{\prime}=\langle 1,\lambda_{k}^{1}=\{\mathcal{O}\},\pi_{k}^{1}=a_{k}\rangle,$ (8)
where $\mathcal{O}$ represents any observation, and $a$ the corresponding primitive action.
Thus, we can combine the primitive actions $c_{k}^{\prime}$ defined in equation (8) and the conventions $c_{k}$ defined in equation (5) to produce the augmented action-convention space $C$ defined as
$C{}$ =$=\{c_{0}^{\prime},c_{1}^{\prime},\dots,c_{|A|-1}^{\prime},c_{|A|}^{\prime},c_{|A|+1},\dots,c_{|A|+K}\}$ (9)
$=\langle\pi_{0}^{1},\pi_{1}^{1},\dots,\pi_{|A|-1}^{1},\pi_{|A|}^{1},\pi_{|A|+1}^{1},\pi_{|A|+1}^{2},\dots,\pi_{|A|+K}^{m_{K}-1},\pi_{|A|+K}^{m_{K}}\rangle,$
with
$|C|=\sum_{j=0}^{K}m_{j}+|A|.$ (10)
This ensures that an agent has access to actions in situations where conventions aren’t applicable or available. Additionally, this increases the learning problem complexity since an agent must learn to effectively utilise conventions as well as primitive actions, given scenarios where both are applicable and available.
To implement action space augmentation with conventions, a few changes must be made to the existing algorithm of interest. In the case of independent Q-learning shown in Algorithm 1, the algorithm will receive an additional input detailing the list of implemented conventions, which in turn gets appended to the existing primitive action space to produce the augmented action-convention space defined in equation (9). The output remains the same, i.e. a policy in the form of an action-value function, however this function will be conditioned on the augmented action-convention space $C$, rather than just the primitive action space $A$. Therefore, we can define Algorithm 2, which acts as the algorithm for independent Q-learning with an augmented action-convention space in a turn based environment.
1:Initialise hyperparameters similar to Algorithm 1
2:Define the list of conventions using $c_{k}=\langle m_{k},\lambda_{k}^{1},\pi_{k}^{1},\lambda_{k}^{2},\pi_{k}^{2},\dots,\lambda_{k}^{m_{k}},\pi_{k}^{m_{k}}\rangle$, with $k=0:K$
3:Translate the primitive actions to conventions using $c_{k}^{\prime}=\langle 1,\{\mathcal{O}\},\pi_{k}^{1}=a_{k}\rangle$, with $k=0:|A|$
4:Combine the primitive actions and conventions to produce the augmented action-convention space $C=\{c_{0}^{\prime},c_{1}^{\prime},\dots,c_{|A|-1}^{\prime},c_{|A|}^{\prime},c_{|A|+1},\dots,c_{|A|+K}\}$
5:Initialise all $Q(O,C)\leftarrow 0$
6:for each episode do
7:     Reset environment and set total time $t\leftarrow 0$
8:     while $S_{t+1}$ is not terminal do
9:         for all players $i$ from 00 to $P-1$ do
10:              Obtain $O^{i}_{t}$ from environment
11:              ${C}_{mask}=\{C|O_{t}^{i}\in\lambda_{k}\}$triangleright Define the action mask using the available and active conventions determined by the set of conditions $\lambda_{k}$ and the current observation $O_{t}^{i}$
12:              ${C}^{i}_{t}=\text{argmax}_{c}(Q^{i}(O^{i}_{t},c)\cap{C}_{mask})$triangleright Choose ${C}^{i}_{t}$ from the available and active conventions based on $O^{i}_{t}$, the action mask ${C}_{mask}$, and the policy derived from $Q^{i}$
13:              $m,k=C[{C}^{i}_{t}]$triangleright Index the augmented action-convention space to determine $m$ and $k$
14:              $A^{i}_{t}=\pi_{k}^{m}(O^{i}_{t})$triangleright Use the convention policy to determine the environment action
15:              Take action $A^{i}_{t}$ in the environment
16:              Receive and store $R_{t+1}$ and $O^{i}_{t+1}$, with $O^{i}_{t+1}\in S_{t+1}$, from environment
17:              $\sum_{f=0}^{P-1}R_{t+1+f}$triangleright Calculate the FAR similar to Algorithm 1
18:              $Q^{i}(O^{i}_{t},{C}^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},{C}^{i}_{t})+\alpha\bigr{[}\sum_{f=0}^{P-1}R_{t+1+f}+\gamma\max_{c}Q^{i}(O^{i}_{t+P},c)-Q^{i}(O^{i}_{t},{C}^{i}_{t})\bigr{]}$
19:              $t\leftarrow t+1$
20:         end for
21:     end while
22:end for
In line 2, we define the list of conventions from $0:K$, followed by the translation of the primitive actions to the special form of single-step conventions $c_{k}^{\prime}$ in line 3. The conventions and primitive actions are then be combined to form the augmented action-convention space $C$, shown in line 4. In line 11, the current observation, produced by the environment for the active player $i$ at time $t$, determines the available and active conventions using the set of conditions $\lambda_{k}$ for each convention. This produces the action mask to indicate which conventions-steps are currently available to the agent, and is often used in deep learning to ensure that an agent choses “legal” actions given a certain state. Notably, this also encapsulates the legal primitive actions given the current observation, since they are contained within the augmented action-convention space. Once the action mask is determined, agent $i$ can start or continue an available or active convention (or take a primitive action) by choosing a convention-step $C_{t}^{i}$ based on the policy derived from $Q^{i}$, the action mask ${C}_{mask}$, and their observation $O^{i}_{t}$, shown in line 12.
In line 13, we use the chosen convention-step $C_{t}^{i}$ to determine the corresponding convention $k$, i.e. the convention which contains the specific convention-step, and its appropriate step $m$, i.e. the step in that convention where the chosen convention-step occurs. The convention $k$ and the step $m$ can then be used to determine the appropriate environment action $A_{t}^{i}$, according to the policy $\pi_{k}^{m}$ and the current observation $O_{t}^{i}$, shown in line 14. If the agent chose a primitive action from the augmented action-conventions, the step $m$ will be equal to one and the convention $k$ will correspond to the primitive action $a_{k}$, according to equation (8). Finally, the environment action $A_{t}^{i}$ is sent to the environment which in turn produces the next observation and reward used to calculate the FAR, similar to Algorithm 1. As is the case with human players, the agents aren’t allowed to communicate which convention is currently being started or followed, but rather only observes the environment action determined by the appropriate convention policy, shown in line 15.
From the environment’s perspective, agents are merely choosing actions based on the observations the environment provided, but in reality the agents are choosing these actions based on a set of predefined convention policies $\pi_{k}$. In the case of deep RL, lines 11 and 13-14 indicate that our approach requires two new layers to be added to the agent’s architecture, namely one before the input layer to determine the available and active conventions, i.e. the action mask, and one at the output to translate the chosen convention-step $C_{t}^{i}$ into an environment action, based on the corresponding convention $k$, the convention’s step $m$, the corresponding policy $\pi_{k}^{m}$, and the current observation $O_{t}$. This is illustrated in Fig. 1, and in the case of deep Q-learning, the action selection step is performed using $\text{argmax}_{c}Q(O_{t},c)$, with the available and active conventions as well as the primitive actions forming the action mask.
Figure 1: Architecture design for a feed forward neural network (NN) with an augmented action-convention space applied. The environment provides the observation $O_{t}$ and receives the environment action $A_{t}$. The size of the input layer is equal to the size of the observation tuple ($\rho$), and each hidden layer has a size equal to the number of atoms for that layer. The size of the output layer is equal to the size of the augmented action-convention space ($|C|$), and the action selection is determined by the MARL algorithm. Finally, the chosen augmented action-convention determines the specific convention $c_{k}$ and its step $m$, which in turn is used to produce the environment action $A_{t}$ according to the policy $\pi_{k}^{m}$ and the observation $O_{t}$.
Agents with an augmented action-convention space can also make use of a shared policy, i.e. each agent will have access to the same network architecture depicted in Fig. 1, as discussed in Section 2.1. Since conventions act as an extension of the existing action space, it can be applied to any MARL method by adding the necessary convention components on top of the existing architecture as well as translating the existing primitive action space into single-step conventions, similar to our discussion on independent Q-learning.
@@ 4 Hanabi
To test the capabilities of conventions, we will use the cooperative Hanabi environment, which was proposed as a frontier for MARL algorithm development by Bard et al. [36]. Hanabi is a 2–5 player card game, best described as a cooperative solitaire [36]. Players have a hand of five cards for two and three player, and four cards for four and five player. Each card has a suit/colour (red, yellow, green, white, or blue) and a rank (1 to 5). The deck comprises 50 cards total, with 10 cards for each suit with a distribution of three 1’s, two 2’s, 3’s, and 4’s, and finally only one 5. The aim of the game is to stack these suits in numerical order from 1 to 5, however players cannot see their own cards (i.e. their hands are hidden), but can see the cards of their fellow players.
@@
Figure 2: An example game of Hanabi as seen from the perspective of player 1. There are four hint tokens left and the players have lost one of their shared life tokens. Player 1 knows about two 4s in their hand and the green, blue and yellow stacks have been partially completed, leading to a current game score of 6/25. It is now player 1’s turn, and they can take the hint (to player 2 or 3), play (from their hand) or discard (from their hand) action.
At the start of the game the players have a shared total of three life tokens and eight hint tokens. If all the hint tokens are depleted, a player cannot take the hint action and must either play or discard a card from their hand. The players are awarded a shared score depending on the number of cards successfully placed on each stack. If the players manage to build the stacks to the maximum of 5 each, the game ends and the players receive a perfect score of 25/25. Alternatively, if the players lose all of their life tokens (called “bombing out”), or if the deck has been depleted, the game also ends. Note, bombing out results in a score of 0/25 independent of how many cards were played successfully, while deck depletion results in a final score equal to the current score. A discard action, or a misplay, will result in the card being removed from the game, and that player must draw a new card from the deck to replenish the missing card (which is also hidden from that player).
@@ 4.1 Human Conventions in Hanabi
Due to the nature of Hanabi, players have too few hint tokens to effectively convey all the needed information (colour and rank) of the 25 playable cards [36]. This becomes even more challenging as the number of players increase and hint tokens become more valuable, as well as the number of playable turns that keep decreasing [^7] As the number of players increase, the cards remaining in the deck decrease, and a single player has less playable turns.. Hints must therefore be used wisely, and must convey information about more than just one card. This is possible since more than one card with a similar rank or suit can be touched (or revealed) by a single hint, along with negative information [^8] Negative information refers to implied information about non-focus cards, e.g. if a player hints that two cards are green, the other cards must therefore be non-green. being given about all the other cards in a hand. Unfortunately, this is not enough, and players who don’t use additional strategies will struggle to reliably achieve a score above 15/25.
@@
Players, therefore, require a way to convey additional information through implicit communication, while still remaining within the rules of the game. As stated by Bard et al. [36]: “This implicit information is not conveyed through the impact that an action has on the environment (i.e., what happens) but through the very fact that another player decided to take this action (i.e., why it happened)”. Players can then reason over the actions of others, and implicitly communicate through them. This is done through the use of conventions, which are built on principles or mutually agreed upon “rules”, that players develop outside the game. For Example:
Conventions Example 1. Let’s assume the players in Fig. 2 have established a convention which states that if an ambiguous hint is given, the focus of that hint is on the left-most card. Therefore, using this convention, player 1 can hint to player 2 that they have two red cards in position one and three respectively (from the left), and player 2 will know (as a result of their convention) that the left-most card must be the playable red 1999This is a common human convention and is based on the “Single card focus” principle as discussed by the H-Group and the implemented conventions found in Appendix A..
Since the game’s release, players have developed their own intricate and extensive conventions, often specific to their group of friends or cooperators, examples include the Board Game Arena [^10] https://forum.boardgamearena.com/viewtopic.php?t=5252 and the H-Group [^11] https://hanabi.github.io/, with the H-Group being the most widely used and adapted.
4.2 Artificial Conventions in Hanabi
To incorporate existing human conventions into MARL, we use conventions as described in Section 3. This process is best described at the hand of an example:
@@ 4.2 Artificial Conventions in Hanabi
Conventions Example 2. In Fig. 2 each player is controlled by an agent, with the current player (referred to as the active player) being agent 1. The convention at hand (as mentioned in Conventions Example 1), which we will refer to as $c_{1}$, states that if an ambiguous hint is given, the focus of that hint is the left-most (or newest) card. Since the condition $\lambda_{1}^{1}$ is met, i.e. player 2 has two cards in their hand that share the same colour or rank and the left-most one is playable, the action {start $c_{1}$} is available for agent 1. Agent 1 chooses to initiate $c_{1}$ which in turn gets translated by the policy $\pi_{1}^{1}$ to an environment action of {hint red to player 2}, ending player 1’s turn. It is now player 2’s turn, and they have just received an ambiguous colour hint from player 1, and given their current observation they can derive that $c_{1}$ is currently active. This triggers the subscribing (and also completing) condition $\lambda_{1}^{2}$ (it is completing since $m_{1}=2$), where agent 2 can now opt in to continue (and complete) $c_{1}$. When agent 2 chooses to subscribe to $c_{1}$, their action is translated by the policy $\pi_{1}^{2}$ to an environment action of {play card in position 1}, subsequently ending player 2’s turn and completing $c_{1}$.
As seen by this example, and discussed in Algorithm 2, the agents are not allowed to convey directly which convention was started or followed, and merely observe the behaviour of the other player through that player’s environment action and their observation. We note that not all MARL scenarios allow for the observation of another agent’s actions, and agents will therefore only make use of their observations when determining the conventions. However, similar to the SMAC environment [42], Hanabi intrinsically allows the monitoring of other agents’ actions, and is included in an agent’s observation vector. Conventions Examples 1 and 2 focused on the “Single card focus” principle, as discussed in Appendix A.2, however there exists more advanced conventions that allow for improved implicit communication. These principles are all aimed at maximising the amount of information given by a single hint, in order to ensure there are enough hint tokens to effectively convey all the needed information of the 25 playable cards.